If x 3 +1/x 3 =2 ,find x+1/x

If x3+1/x3=2,find x+1/x

Grade:10

2 Answers

Karthik
9 years ago
x+1/x = 2/(x2+(1/x2)-1)
sowmya
9 years ago
given x3+1/x3=2
then(x+1/x) (x2+1/x2-1) =2
(x+1/x) ((x+1/x)2-3) =2
let x+1/x=t
then (t) (t2-3) =2
which is satisfied when t=2
hence t=2 i.e x+1/x=2